OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a_1 = 4 and a_6 = 34 ? 182 186 190 194

OpenStudy (alexwee123):

didn't i just solve a question like this a minute ago?

OpenStudy (anonymous):

similar yes

OpenStudy (anonymous):

A formula you should know: a_n = a_1 + (n-1)d where d is the common difference. You are given a_1 = 4 and a_6 = 34. This will enable you to determine d. 34 = a_6 = a_1 + 5d = 4 + 5d 30 = 5d 6 = d Therefore, a_n = 4 + (n - 1)6 = 4 + 6n - 6 = 6n - 2. a_32 = 6*32 - 2 = 190.

OpenStudy (anonymous):

The answers 190